City A and VCity B are 440 miles apart. If Jack leaves city A traveling 50 mph and Jill leaves city B traveling 60 mph, how long will they travel if they drive toward each other?
Answer by ptaylor(1894)

Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
I'm assuming that the question is "how long will they travel before they meet"
Let t=amount of time they travel before they meet
Distance Jack travels=50t
Distance Jill travels=60t
Now we know that when the above two distances add up to 440 mi, Jack and Jill will have met, so:
50t+60t=440
110t=440 divide both sides by 110
t=4 hrs--------------------------amount of time they travel before they meet.
CK
50*4+60*4=440
200+240=440
440=440

Hope this helps---ptaylor